OpenStudy (avountae2003):

I have these on my study guide but for some reason I keep getting them wrong can someone break it down for me. 1. 2+4(x-7)=7-5(x-4) 2. 2(5z-1)-2(z+9)=3(z+1) This one is a fraction: -5/9x+2/3=5/9

OpenStudy (anonymous):

can you show how you did these?

OpenStudy (avountae2003):

1. 2+4x-28=7-5x+20 4x-26=27+5x +26 4x=53+5x -5x 9x=53 /9 53/9=x

OpenStudy (avountae2003):

2. 10z-2-2z-18=3z+3 8z-20=3z+3 +20 8z=3z+23 -3 5z=23 /5 23/5=z

OpenStudy (avountae2003):

Thats all I know because the fraction one I dont understand at all.

OpenStudy (anonymous):

ok. give me a sec

OpenStudy (avountae2003):

ok

OpenStudy (mathstudent55):

1. 2+4x-28=7-5x+20 4x-26=27+5x ^ | how did this become +? It was - in the line abnove.

OpenStudy (mathstudent55):

^above

OpenStudy (avountae2003):

oh I accidentally changed it

OpenStudy (mathstudent55):

Interestingly, later you made another mistake, and the answer came out correct.

OpenStudy (anonymous):

I ended up with the same thing you did for both.

OpenStudy (mathstudent55):

1. 2 + 4x - 28 = 7 - 5x + 20 4x - 26 = 27 - 5x +5x +5x ------------------ 9x - 26 = 27 + 26 +26 -------------- 9x = 53 x = 53/9

OpenStudy (avountae2003):

When I put them in my study it guide it say I have done it wrong.

OpenStudy (mathstudent55):

Your final answer to problem 1 is correct, but you made two mistakes which somehow canceled each other.

OpenStudy (avountae2003):

okay could you show me what I did wrong

OpenStudy (mathstudent55):

Problem 2. 2(5z-1)-2(z+9)=3(z+1) 10z-2-2z-18=3z+3 <---- correct 8z-20=3z+3 <----- correct +20 --------------- 8z=3z+23 <------- correct -3 <----- should read -3z 5z=23 /5 23/5=z <----- correct final answer

OpenStudy (mathstudent55):

You did problem 2 well.

OpenStudy (avountae2003):

Thank you but for some reason it is wrong

OpenStudy (anonymous):

For the third one is it like this with the x beside -5/9? just want to make sure it's not in the denominator \[-\frac{ 5 }{ 9 }x+\frac{ 2 }{ 3 }=\frac{ 5 }{ 9}\]

OpenStudy (mathstudent55):

The answer is correct. Do they ask for a decimal instead of a fraction?

OpenStudy (avountae2003):

Yes peachpi and it ask for a integer or simplified fraction

OpenStudy (anonymous):

ok. you can do the same thing you did for the others. Start by subtracting 2/3 from both sides, then you'll need to get a common denominator to subtract the fractions.